OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point shown and is parallel to the line shown. line through the points 0 comma 6 and 10 comma 0 with a point 4 comma negative 6 not on the line

OpenStudy (anonymous):

y = -2x + 2 y = 2x - 14 y = one halfx - 8 y = -one halfx - 4

OpenStudy (anonymous):

ok so wait it would look like this? (0,6) (10,0) (4,-6)?

OpenStudy (anonymous):

do you know what the slope what the slope-intercept formula is?

OpenStudy (anonymous):

yes isn't y=mx+b

OpenStudy (anonymous):

plug your numbers in that are the 2 sets of coordinates on the number line